# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ 6.3. Planar open set with non-locally separating simple closed curve
Let $X$ be a Peano continuum, $\gamma$ be a simple closed curve on $X$, and $\mathcal{S}$ be a brick partition of $X$. A spaced circular path $c=(c(0),\dots,c(\ell-1),c(0))$ on $\mathcal{S}$ is $\gamma$-compatible if
@@
Let $X$ be a Peano continuum with no locally separating points and let $\gamma$ be a non-locally separating simple closed curve in an open planar subset $O\subseteq X$. For each brick partition $\mathcal{U}$ of $X$, there are a brick partition $\mathcal{V}\preceq\mathcal{U}$ and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}$.
@@
Let $\varepsilon$ be less than half of $\min\{\operatorname{diam}(U):U\in\mathcal{U}\}$ and let $\delta$ be the uniform local connectedness constant of $U$ relative to $\varepsilon$. We prove that if two points $y_{0},y_{1}\in\gamma\cap\operatorname{Cl}(U)$ have distance less than $\delta$, then they belong to the same connected component. Indeed, in this case there is an arc $\lambda$ from $y_{0}$ to $y_{1}$ in $U\cup\{y_{0},y_{1}\}$ of diameter less than $\varepsilon$. The arc $\lambda$ separates $O$ in two regions, one of which meets $U$, has diameter less that $2\varepsilon$, and contains in its interior an open interval $I\subseteq\gamma$ with endpoints $y_{0},y_{1}$. But then such region is too small to contain any $U^{\prime}\in\mathcal{U}$. Moreover, it cannot even meet any $U^{\prime}\neq U$, since $\lambda$ would separate $U^{\prime}$ and $\lambda$ is contained in $\operatorname{Cl}(U)$. Therefore it has to be contained in $\operatorname{Cl}(U)$. But then $I\subseteq U$, and thus $y_{0},y_{1}$ are in the same connected component of $\gamma\cap\operatorname{Cl}(U)$. This shows that there are only finitely many components. It also shows that $\gamma\cap\operatorname{Bd}(U)$ is totally disconnected, since the open interval between any two points in the same connected component of $\gamma\cap\operatorname{Cl}(U)$ is contained in $U$.
Suppose towards a contradiction that $\{x\}$ is a trivial connected component of $\operatorname{Cl}(U)\cap\gamma$, so in particular $x\in\operatorname{Bd}(U)$. Let $\delta_{0}$ be the minimum between $\delta$ and the uniform local connectedness constant of $\bigvee\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$ relative to $\varepsilon$. Take two points $y_{0},y_{1}\in\gamma\cap\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$ whose distance is smaller than $\delta_{0}$ and such that $x\in\gamma(y_{0},y_{1})$. Then there is an arc $\lambda$ of diameter less than $\varepsilon$ connecting $y_{0},y_{1}$ in $\bigvee\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$. In particular $\lambda\subseteq O$, so it separates $O$ in two regions, one of which contains $x$ in its interior and has diameter less that $2\varepsilon$. Arguing as in the first part of the proof we conclude that $U$ cannot be contained in such region, so $x\not\in\operatorname{Cl}(U)$, a contradiction. ∎
@@
Let $X$ be a Peano continuum with no locally separating points and let $\gamma$ be a non-locally separating simple closed curve contained in an open planar subset $O\subseteq X$. For each brick partition $\mathcal{V}$ of $X$ and path $v$ on $\mathcal{V}$, there are a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$, and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}^{\prime}$ disjoint from $v^{\prime}$.
@@
For each $11<i<\operatorname{lh}(v)-$ 1, let $x_{i}$ be a point in $\operatorname{Cl}(v(i-1))\cap\operatorname{Cl}(v(i))\cap(v(i-1)\vee v(i))$ which does not belong to $\gamma$, which exist by 4.2.(3). Let $\lambda_{i}$ be an arc in $v(i)\cup\{x_{i-1},x_{i}\}$ connecting $x_{i-1}$ to $x_{i}$ and disjoint from $\gamma$, which exist by 4.3.(4).
@@
As in Theorem 6.6, it is sufficient to find a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ and a circular path $c$ disjoint form $v^{\prime}$ satisfying the conditions in Definition 6.3.
Let $O\subseteq X$ be a planar open set and $\gamma\subseteq O$ a simple closed curve which is not locally separating. Fix any brick partition $\mathcal{W}$ of $X$ which is fine enough that $\lvert\operatorname{Star}(\gamma,\mathcal{W})\rvert\geq 4$. Let $\mathcal{S}\preceq\mathcal{W}$ be a brick partition and let $C$ be $\gamma$-compatible spaced circular path $C$ on $\mathcal{S}$, using Lemma 6.8. By the assumption on $\mathcal{W}$, the length $\ell$ of $C$ is at least $4$.
@@
Suppose a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. Up to a refinement, we can suppose that $v$ is a path, by (Upgrading Walks to Paths). (Theorem 6.2). By Proposition 6.9, there are a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}^{\prime}$ which is disjoint from $v^{\prime}$.
Now $\rho^{\mathcal{V}^{\prime}}_{\mathcal{U}}\cdot c$ is confined in $D$, by point 1 of Definition 6.7, so (6.1) follows from (6.13). Since changing the starting vertex of $C$ does not affect the winding number, and flipping the direction of $C$ just flips the sign of $\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c)$, we can assume that $c(0)\in\mathcal{V}^{\prime}(C(0))$ and that the first time $c$ exits $\mathcal{V}^{\prime}(C(0))$ it enters $\mathcal{V}^{\prime}(C(1))$. Item (2) in Definition 6.7 then implies that both $c$ and $C$ induce brick partitions of $\gamma$, with the former refining the latter. It follows that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c$ monotonically refines $C$, thus by Lemma 3.2 we can conclude that $\lvert\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c)\rvert=\ell$. ∎
@@ 6.4. Circular covering
We recall from the introduction that a Peano continuum $X$ has a circular covering if it admits a covering $O_{0},\dots,O_{\ell-1}$ in $\ell\geq 4$ connected open subsets such that $\operatorname{Cl}(O_{i})\cap\operatorname{Cl}(O_{j})\neq\emptyset$ if and only if $\lvert i-j\rvert\leq 1\text{ mod}\,\ell$, and $O_{i}\setminus\bigcup_{j\neq i}O_{j}$ is connected for each $i<\ell$.
@@
$C_{i}\coloneqq\operatorname{Cl}(O_{i})\setminus\bigcup_{j\neq i}O_{j}.$
These sets are pairwise disjoint, since $x\in C_{j}\cap C_{k}$ with $j\not=k$ if and only if $x\not\in O_{i}$ for every $i<\ell$, which is not possible.
@@
(6.14) $\operatorname{Star}(C_{i},\mathcal{U})\cap\operatorname{Star}(C_{j},\mathcal{U})=\emptyset\text{ for each }i\neq j,$
@@
(6.15) $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})\cap\operatorname{Star}(O_{j}\cap O_{j+1\text{ mod}\,\ell},\mathcal{U})=\emptyset\text{ for each }i\neq j.$
Since $C_{i}$ is connected, $\mathcal{U}_{i}=\operatorname{Star}(C_{i},\mathcal{U})$ is $E$-connected. Given $i<\ell$, divide $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})$ into two sets $\mathcal{P}_{i},\mathcal{Q}_{i+1\text{ mod}\,\ell}$ based on whether $U$ is such that
@@
Using (6.14) and (6.15) we see that if $U\in\mathcal{P}_{i}$ then all the elements of any of the shortest paths connecting it to $\mathcal{U}_{i}$ are either in $\mathcal{U}_{i}$, or in $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})$, and therefore in $\mathcal{P}_{i}$. A similar observation holds for $\mathcal{Q}_{i}$.
We can thus conclude that the sets $\mathcal{Q}_{i}\cup\mathcal{U}_{i}\cup\mathcal{P}_{i}$ are $E$-connected and form a partition of $\mathcal{U}$. The amalgam $\mathcal{S}$ of $\mathcal{U}$ obtained from such partition is the desired cycle of $\ell$ elements. ∎
@@
By Proposition 6.11, there is a brick partition $\mathcal{S}$ of $X$ which is a cycle of length $\ell\geq 4$. Let $C$ be a circular path around $\mathcal{S}$ (in particular, $\operatorname{lh}(C)=\ell$). By (Upgrading Walks to Paths). (Theorem 6.2), there are a brick partition $\mathcal{U}\preceq\mathcal{S}$ and a spaced path $u$ such that $\rho^{\mathcal{U}}_{\mathcal{S}}\cdot u$ monotonically refines $C{}^{\frown}C$, which is possible since $C{}^{\frown}C$ is uncrossed. In particular, $\operatorname{wd}_{C}(\rho^{\mathcal{U}}_{\mathcal{S}}\cdot u)=2\ell$, by Lemma 3.2.
Suppose a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. Suppose without loss on generality that $v(0)\in\mathcal{V}(C(0))$. Note that for any initial segment $t$ of $v$ such that $\rho^{\mathcal{V}}_{\mathcal{S}}(t(-1))\in C(k)$ we get
@@
Let $s_{i}$ be an initial segment of $v_{i}$ such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot s_{i}$ monotonically refines $t$. Then, again by Lemma 3.2,
$\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot s_{i})=\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot t)>\ell.$
To lighten notation, let us refer to $\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{U}}\cdot r)$ as the winding number of $r$, for any walk $r$ on $\mathcal{V}^{\prime}$.
Let us partition the set of the $V$’s in $\mathcal{V}^{\prime}(C(k_{0}))$ which belong to $v_{0}\cup v_{1}$ into two sets $A,B$, depending on whether the winding number of the corresponding initial segment up to $V$ is smaller or greater than $\ell$, in absolute value (by (6.16) the value $\ell$ is never achieved). Both of these sets are nonempty, as witnessed by the $s_{i}$’s. Let $z$ be a path from $A$ to $B$ in $\mathcal{V}^{\prime}(C(k_{0}))$. There are two cases.
Case 1: There is $i<2$ such that both endpoints of $z$ are contained in $v_{i}$, and we might as well assume $i=1$. Let $v^{\prime}{}^{\frown}c$ be the lasso path which follows $v_{0}$ from beginning to end, then $v_{1}$ backwards until it hits $z$ for the first time (this is $v^{\prime}$), then follows $z$ (possibly backwards), then $v_{1}$ forwards until it hits $z$ again (this is $c$). Since $z$ is included in $\mathcal{V}^{\prime}(C(k_{0}))$ it does not contribute to the winding number of $c$, which is therefore equal to the difference between the winding numbers of the initial segments of $v_{1}$ at either end of $z$, which are equal modulo $\ell$ by (6.16). But these are in $A$ and $B$, respectively, so the absolute value of the difference is greater or equal than $\ell$, and (6.2), is satisfied.
Case 2: Otherwise, let $v^{\prime}{}^{\frown}c$ be the lasso path which follows $v_{0}$ from the beginning until it hits $z$ (this is $v^{\prime}$), then continues following $v_{0}$ until the end, then $v_{1}$ backwards until it hits $z$, then follows $z$, possibly backwards (this is $c$). As above, $z$ does not contribute to the winding number of $c$, which is therefore equal to the difference between the winding numbers of the final segments of $v_{1}$ and $v_{0}$ at either end of $z$. But since $v_{0},v_{1}$ have the same winding number by (6.17), the difference between the winding numbers of the final segments is equal to the difference of the winding numbers of the respective initial segments. We therefore conclude as in Case 1.
@@
A continuum is unicoherent if it cannot be written as the union of two subcontinua whose intersection is disconnected. If $X$ has a circular covering covering $O_{0},\dots,O_{\ell-1}$, it is not unicoherent, as witnessed by $\operatorname{Cl}(O_{0})$ and $\bigcup_{i\neq 0}\operatorname{Cl}(O_{i})$.
@@
On the other hand, it is clear that the torus and the Klein bottle have circular coverings. We conclude by showing that the connected sum of a surface $X$ which has a circular covering, with any other surface $X^{\prime}$, has a circular covering: suppose $O_{0},\dots,O_{\ell-1}$ is a circular covering of $X$; take an open disk $D$ in $O_{0}\setminus\bigcup_{j\neq 0}O_{j}$, and any open disk $D^{\prime}$ in $X^{\prime}$, and glue $X,X^{\prime}$ along $\operatorname{Bd}(D),\operatorname{Bd}(D^{\prime})$. Then $(O_{0}\setminus D)\cup(X^{\prime}\setminus D^{\prime}),O_{1},\dots,O_{\ell-1}$ is a circular covering of $X\#X^{\prime}$. ∎

### v3

@@ 6.4. Planar open set with non-locally-separating simple closed curve
Let $X$ be a Peano continuum, $\gamma$ be a simple closed curve in $X$, and $\mathcal{S}$ be a brick partition of $X$. A spaced circular path $c=\langle c(0),\dots,c(\ell-1),c(0)\rangle$ on $\mathcal{S}$ is $\gamma$-compatible if
@@
Throughout this subsection we make repeated use of 4.2.(3) and 4.3.(4) applied to a non-locally-separating simple closed curve $\gamma$ in a Peano continuum with no locally separating points. Indeed, this is possible the since simple closed curves in Peano continua with no locally separating points necessarily have empty interior.
@@
Let $X$ be a Peano continuum with no locally separating points and let $\gamma$ be a non-locally-separating simple closed curve in an open planar subset $O\subseteq X$. For each brick partition $\mathcal{U}$ of $X$, there are a brick partition $\mathcal{V}\preceq\mathcal{U}$ and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}$.
@@
Let $\varepsilon$ be less than half of $\min\{\operatorname{diam}(U):U\in\mathcal{U}\}$, fix once and for all some $U\in\mathcal{U}$, and let $\delta$ be the uniform local connectedness constant of $U$ relative to $\varepsilon$. We prove that if two points $y_{0},y_{1}\in\gamma\cap\operatorname{Cl}(U)$ have distance less than $\delta$, then they belong to the same connected component. Indeed, in this case there is an arc $\lambda$ from $y_{0}$ to $y_{1}$ in $(U\setminus\gamma)\cup\{y_{0},y_{1}\}$ of diameter less or equal than $\varepsilon$. The arc $\lambda$ separates $O$ in at least two regions, one of which meets $U$, has diameter less that $2\varepsilon$, and contains in its interior an open interval $I\subseteq\gamma$ with endpoints $y_{0},y_{1}$. But then such region is too small to contain any $U^{\prime}\in\mathcal{U}$. Moreover, it cannot even meet any $U^{\prime}\neq U$, since $\lambda$ would separate $U^{\prime}$ and $\lambda$ is contained in $\operatorname{Cl}(U)$. Therefore it has to be contained in $\operatorname{Cl}(U)$. But then $I\subseteq U$, and thus $y_{0},y_{1}$ are in the same connected component of $\gamma\cap\operatorname{Cl}(U)$. This shows that there are only finitely many components. It also shows that $\gamma\cap\operatorname{Bd}(U)$ is totally disconnected, since the open interval between any two points in the same connected component of $\gamma\cap\operatorname{Cl}(U)$ is contained in $U$.
Suppose towards a contradiction that $\{x\}$ is a trivial connected component of $\operatorname{Cl}(U)\cap\gamma$, so in particular $x\in\operatorname{Bd}(U)$. Let $\delta_{0}$ be the minimum between $\delta$ and the uniform local connectedness constant of $\bigvee\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$ relative to $\varepsilon$. Take two points $y_{0},y_{1}\in\gamma\cap\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$ whose distance is smaller than $\delta_{0}$ and such that $x\in\gamma(y_{0},y_{1})$. Then there is an arc $\lambda$ of diameter less than $\varepsilon$ connecting $y_{0},y_{1}$ in $\bigvee\operatorname{Star}(x,\mathcal{U}\setminus\{U\})$. In particular $\lambda\subseteq O$, so it separates $O$ in at least two regions, one of which contains $x$ in its interior and has diameter less that $2\varepsilon$. Arguing as in the first part of the proof we conclude that $U$ cannot be contained in such region, so $x\not\in\operatorname{Cl}(U)$, a contradiction. ∎
@@
Let $X$ be a Peano continuum with no locally separating points and let $\gamma$ be a non-locally-separating simple closed curve contained in an open planar subset $O\subseteq X$. For each brick partition $\mathcal{V}$ of $X$ and path $v$ on $\mathcal{V}$, there are a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$, and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}^{\prime}$ disjoint from $v^{\prime}$.
@@
For each $11\leq i<\operatorname{lh}(v)-$ 1, let $x_{i}$ be a point in $\operatorname{Cl}(v(i-1))\cap\operatorname{Cl}(v(i))\cap(v(i-1)\vee v(i))$ which does not belong to $\gamma$. Let $\lambda_{i}$ be an arc in $v(i)\cup\{x_{i-1},x_{i}\}$ connecting $x_{i-1}$ to $x_{i}$ and disjoint from $\gamma$.
@@
Let $O\subseteq X$ be a planar open set and $\gamma\subseteq O$ a simple closed curve which is not locally separating. Fix any brick partition $\mathcal{W}$ of $X$ which is fine enough that $\lvert\operatorname{Star}(\gamma,\mathcal{W})\rvert\geq 4$. Let $\mathcal{S}\preceq\mathcal{W}$ be a brick partition and let $C$ be $\gamma$-compatible spaced circular path $C$ on $\mathcal{S}$, using Lemma 6.8. By the assumption on $\mathcal{W}$, the circular length $\ell$ of $C$ is at least $4$.
@@
Suppose a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. As in Theorem 6.6, it is sufficient to find a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ and a circular path $c$ disjoint from $v^{\prime}$ satisfying the conditions in Definition 6.3. Up to a refinement, we can suppose that $v$ is a path, by (Upgrading Walks to Paths). (Theorem 6.2). By Proposition 6.9, there are a brick partition and a path $(\mathcal{V}^{\prime},v^{\prime})\preceq(\mathcal{V},v)$ and a $\gamma$-compatible spaced circular path $c$ on $\mathcal{V}^{\prime}$ which is disjoint from $v^{\prime}$.
Now $\rho^{\mathcal{V}^{\prime}}_{\mathcal{U}}\cdot c$ is confined in $D$, by point (1) of Definition 6.7, so (6.1) follows from (6.13). Since changing the starting vertex of $C$ does not affect the winding number, and flipping the direction of $C$ just flips the sign of $\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c)$, we can assume that $c(0)\in\mathcal{V}^{\prime}(C(0))$ and that the first time $c$ exits $\mathcal{V}^{\prime}(C(0))$ it enters $\mathcal{V}^{\prime}(C(1))$. Item (2) in Definition 6.7 then implies that both $c$ and $C$ induce brick partitions of $\gamma$, with the former refining the latter. It follows that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c$ monotonically refines $C$, thus by Lemma 3.3 we can conclude that $\lvert\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot c)\rvert=\ell$. ∎
@@ 6.5. Circular covering
We recall from the introduction that a Peano continuum $X$ has a circular covering if it admits a covering $O_{0},\dots,O_{\ell-1}$ in $\ell\geq 4$ connected open subsets such that $\operatorname{Cl}(O_{i})\cap\operatorname{Cl}(O_{j})\neq\emptyset$ if and only if $\lvert i-j\rvert\leq 1\text{ mod}\,\ell$, and $O_{i}\setminus\bigcup_{j\neq i}O_{j}$ is connected for each $i<\ell$. Examples of Peano continua with no locally separating points admitting a circular covering are the torus, the Menger curve or the Sierpiński carpet.
@@
$C_{i}\coloneqq O_{i}\setminus\bigcup_{j\neq i}O_{j}.$
These sets are closed and pairwise disjoint, since $x\in C_{j}\cap C_{k}$ with $j\not=k$ if and only if $x\not\in O_{i}$ for every $i<\ell$, which is not possible.
@@
(6.14) $\operatorname{Star}(C_{i},\mathcal{U})\cap\operatorname{Star}(C_{j},\mathcal{U})=\emptyset,\text{ for each }i\neq j,$
@@
(6.15) $\begin{gathered}d_{\mathcal{U}}\left(\operatorname{Star}\left(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U}\right),\operatorname{Star}\left(O_{j}\cap O_{j+1\text{ mod}\,\ell},\mathcal{U}\right)\right)\geq 2,\text{ for each }i\neq j,\\ d_{\mathcal{U}}\left(\operatorname{Star}\left(O_{i}\cap O_{i+1\bmod\ell},\mathcal{U}\right),\operatorname{Star}(C_{j},\mathcal{U})\right)\geq 2,\text{ for all }j\notin\{i,i+1\text{ mod}\,\ell\}.\end{gathered}$
Since $C_{i}$ is connected, $\mathcal{U}_{i}\coloneqq\operatorname{Star}(C_{i},\mathcal{U})$ is $E$-connected, and all $\mathcal{U}_{i}$’s are pairwise disjoint by (6.14). Given $i<\ell$, divide $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})$ into two sets $\mathcal{P}_{i},\mathcal{Q}_{i+1\text{ mod}\,\ell}$ based on whether $U$ is such that
@@
Notice that, by (6.15), if $U\in\mathcal{P}_{i}$ then all the elements of any path of minimal length, call it $P$, connecting $U$ to $\mathcal{U}_{i}$ are either in $\mathcal{U}_{i}$, or in $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})$, and therefore in $\mathcal{P}_{i}$. Indeed, if $U^{\prime}$ is the first step in $P$ that does not belong to $\operatorname{Star}(O_{i}\cap O_{i+1\text{ mod}\,\ell},\mathcal{U})$, then by (6.15) we have $U^{\prime}\cap O_{j}=\emptyset$ for all $j\neq i,i+1\text{ mod}\,\ell$. It thus follows that $U^{\prime}$ has to belong to either $\mathcal{U}_{i}$ or $\mathcal{U}_{i+1\text{ mod}\,\ell}$ (but not to both, by (6.14)), and the latter is not possible as otherwise $U\in\mathcal{Q}_{i}$.
A similar observation holds for $\mathcal{Q}_{i}$. We can thus conclude that the sets $\mathcal{Q}_{i}\cup\mathcal{U}_{i}\cup\mathcal{P}_{i}$ are $E$-connected and form a partition of $\mathcal{U}$. The amalgam $\mathcal{S}$ of $\mathcal{U}$ obtained from such partition is the desired cycle of $\ell$ elements. ∎
@@
By Proposition 6.11, there is a brick partition $\mathcal{S}$ of $X$ whose nerve graph is a cycle of length $\ell\geq 4$. Let $C$ be a circular path around $\mathcal{S}$ (in particular, the circular length of $C$ is $\ell$). By (Upgrading Walks to Paths). (Theorem 6.2), there are a brick partition $\mathcal{U}\preceq\mathcal{S}$ and a spaced path $u$ such that $\rho^{\mathcal{U}}_{\mathcal{S}}\cdot u$ monotonically refines $C{}^{\frown}C$, which is possible since $C{}^{\frown}C$ is uncrossed. In particular, $\operatorname{wd}_{C}(\rho^{\mathcal{U}}_{\mathcal{S}}\cdot u)=2\ell$, by Lemma 3.3.
Suppose a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ are given. Suppose without loss of generality that $v(0)\in\mathcal{V}(C(0))$. Note that for any initial segment $t$ of $v$ such that $\rho^{\mathcal{V}}_{\mathcal{S}}(t(-1))\in C(k)$ we get
@@
Let $s_{i}$ be an initial segment of $v_{i}$ such that $\rho^{\mathcal{V}^{\prime}}_{\mathcal{V}}\cdot s_{i}$ monotonically refines $t_{0}$. Then, again by Lemma 3.3,
$\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot s_{i})=\operatorname{wd}_{C}(\rho^{\mathcal{V}}_{\mathcal{S}}\cdot t_{0})>\ell.$
To lighten notation, let us refer to $\operatorname{wd}_{C}(\rho^{\mathcal{V}^{\prime}}_{\mathcal{S}}\cdot r)$ as $\omega_{r}$, for any walk $r$ on $\mathcal{V}^{\prime}$.
Let us partition the set of the $V$’s in $\mathcal{V}^{\prime}(C(k_{0}))$ which belong to $v_{0}\cup v_{1}$ into two sets $A,B$, depending on whether $|\omega_{r}|>\ell$ or $|\omega_{r}|<\ell$, where $r$ is the initial segment of either $v_{0}$ or $v_{1}$ ending in $V$ (by (6.16) the value $\ell$ is never achieved). Both of these sets are nonempty, as witnessed by the $s_{i}$’s. Let $z$ be a path from $A$ to $B$ in $\mathcal{V}^{\prime}(C(k_{0}))$. There are two cases.
Case 1: There is $i<2$ such that both endpoints of $z$ are contained in $v_{i}$, and we might as well assume $i=1$. Let $v^{\prime}{}^{\frown}c$ be the lasso path in $\mathcal{V}^{\prime}$ which follows $v_{0}$ from beginning to end, then $v_{1}$ backwards until it hits $z$ for the first time (this is $v^{\prime}$), then follows $z$ (possibly backwards), then $v_{1}$ forwards until it hits $z$ again (this is $c$). Since $z$ is included in $\mathcal{V}^{\prime}(C(k_{0}))$ it does not contribute to the winding number of $c$, which is therefore equal to the difference $\omega_{r_{0}}-\omega_{r_{1}}$ where $r_{0},r_{1}$ are the initial segments of $v_{1}$ at either end of $z$. By (6.16) we have $\omega_{r_{0}}\equiv\omega_{r_{1}}\text{ mod}\,\ell$. Furthermore, we have that either $r_{0}\in A$ and $r_{1}\in B$ or viceversa, hence $|\omega_{c}|=|\omega_{r_{0}}-\omega_{r_{1}}|\geq\ell$, and (6.2), is satisfied.
Case 2: Otherwise, let $v^{\prime}{}^{\frown}c$ be the lasso path which follows $v_{0}$ from the beginning until it hits $z$ (this is $v^{\prime}$), then continues following $v_{0}$ until the end, then $v_{1}$ backwards until it hits $z$, then follows $z$, possibly backwards (this is $c$). As above, $z$ does not contribute to the winding number of $c$, which is therefore equal to the difference $\omega_{r_{0}}-\omega_{r_{1}}$, where $r_{0},r_{1}$ are the final segments of $v_{0}$ and $v_{1}$ at either end of $z$. But since $v_{0},v_{1}$ have the same winding number by (6.17), $\omega_{r_{0}}-\omega_{r_{1}}$ is equal to the difference of the winding numbers of the respective initial segments. We therefore conclude as in Case 1.
@@
We close this section showing that the sphere and the real projective plane are the only two closed surfaces without a circular covering.
A continuum is unicoherent if it cannot be written as the union of two subcontinua whose intersection is disconnected. If $X$ has a circular covering $O_{0},\dots,O_{\ell-1}$, it is not unicoherent, as witnessed by $\operatorname{Cl}(O_{0})$ and $\bigcup_{i\neq 0}\operatorname{Cl}(O_{i})$.
@@
On the other hand, it is clear that the torus and the Klein bottle have circular coverings. Recall that, by the classification theorem for compact surfaces [MR2766102]*Theorem 6.15 every compact surface is either the sphere, a connected sum of copies of the torus, or a connected sum of copies of the real projective plane $\mathbb{R}P^{2}$. Since $\mathbb{R}P^{2}\#\mathbb{R}P^{2}$ is homeomorphic to the Klein bottle, we conclude by showing that the connected sum of a surface $X$ which has a circular covering, with any other surface $X^{\prime}$, has a circular covering: suppose $O_{0},\dots,O_{\ell-1}$ is a circular covering of $X$; take an open disk $D$ in $O_{0}\setminus\bigcup_{j\neq 0}O_{j}$, and any open disk $D^{\prime}$ in $X^{\prime}$, and glue $X,X^{\prime}$ along $\operatorname{Bd}(D),\operatorname{Bd}(D^{\prime})$. Then $(O_{0}\setminus D)\cup(X^{\prime}\setminus D^{\prime}),O_{1},\dots,O_{\ell-1}$ is a circular covering of $X\#X^{\prime}$. ∎